10) You have a 55 year old male in LTC facility that has a PEG tube. Height is 6 ‘2and weight is 200 lbs. The patient is getting enteral feeding. You perform all the necessary calculations and determine his BEE is 1884 Kcal. You do an adjustment to the BEE of 1.2. You do an adjustment to the Protein requirements of 1.3. You decide to use Boost Plus which is 1.52 Kcal /ml. It has 78% free water in each 237ml can. Each can has 16% protein. How many cans a day will be required to meet the Adjusted BEE?? (Round to the nearest bottle) If this was an enteral pump how fast would you set the pump in mVh to deliver this in 24 hour period? Does the protein meet the requirement???
Answer #1

Patient BEE is 1884kcal

Each bottle has 237ml of boost..

Cans required to meet the adjusted BEE is 1884/237= 7.9=8bottles

To calculate the ml / hr

Multiply 237by 8 and divide by 60minutes

237X8 = 1,896/60 = 31.6 = 32ml/hr or 768ml/ 24hours

Protein requirements

Patient weight 200ibs or 91kg

91kg/1.5g/day

136g/day

1g protein = 4kcal energy

136g X4

544 kcal/day from protein..